# Towards SLO-Optimized LLM Serving via Automatic Inference Engine Tuning

### v1

@@ IV-G Random Forest-based POF Learning
Figure 6: Probability density function (PDF) of the request input and output lengths for the four request traces used in experiments.
IV-H Ensuring SLO Stability
@@
Since the order in which requests arrive can affect the performance metrics obtained by stress testing, the SLOs of performance tuning vary with different request arrival orders. Through extensive experiments, we find that although the optimal objective value varies with the requests’ arrival order, the corresponding parameter configurations are often very similar. Consequently, after completing the tuning process, we re-conduct the stress testing multiple times with different request arrival orders and select the worst-case objective value as the SLO to ensure the stability of the SLOs. Finally, we summarize the overall tuning procedure of SCOOT in Algorithm 1.
V Experimental Evaluation
V-A Experiment Setup
Implementation. We implement SCOOT with Python 3.8. Pytorch [26] and GPytorch [27] are used to implement the Sobol sequence-based Quasi-Monte Carlo sampling and GP, respectively. For configuration suggestion, we implement the solver with pymoo [28]. Moreover, the random forest regression is implemented using sklearn [29].
@@ V-A Experiment Setup
Computing Resources. Experiments are conducted on our testbed. Multiple sets of computing resources are utilized, including 2A10, 4A10, 2A100, and 4A100, where A10 represents the NVIDIA A10 24GB GPU, A100 denotes the NVIDIA A100 80GB GPU, and the integer before the GPU type indicate the number of GPUs. For each set of computing resources, 256GB CPU memory and one 2.90GHz 32-core Intel(R) Xeon(R) CPU are equipped. Besides, A100 GPUs are connected over NVLink, and A10 GPUs are connected over PCIE.
Request Trace. We collect request traces from four LLM inference services at AntGroup, which include four applications of text-to-sql (SQL), chatbot (BOT), classification (CLS), and recommendation (REC). SQL and BOT are interactive applications, CLS is an online application while REC is an offline application. The distribution of request input lengths and output lengths for the four LLM-empowered applications are presented in Fig. 6.
Inference Engine. We adopt vLLM-v0.4.2 as the inference engine, whose default configuration of to-be-tuned parameters is the same as the latest version.
LLM. We conducted experiments with the LLAMA2-7B and LLAMA2-13B models [18] because the model architecture of LLAMA is representative. Besides, in practice, 7B and 13B are common parameter sizes, and fine-tuned 7B and 13B LLMs can fulfill the requirements for most applications.
@@
Workflow. In each configuration evaluation, when the inference engine is started with the suggested configuration, we send requests from the request trace for 100 seconds and the request arrival times are generated using Poisson distribution with various request rates. For different request traces and GPU types, the request rate is different. For A10, the request rate for SQL, BOT, CLS, and REC are 5,5, 10, and 15, respectively. For A100, the request rates of these request traces are twice that of A10. The request rate is set according to the actual workload of the applications as well as considering the computing capability and GPU memory capacity of GPUs.
Metrics. Since different applications aim to optimize different performance metrics. For SQL and BOT, TTFT and TPOT are leveraged as optimization objectives, while for CLS and REC, 95th percentile latency and request throughput are utilized as performance metrics, respectively. In the experimental results, we report the improvement or reduction of these metrics over the default configuration.
@@
(c) Avg TTFT and TPOT reduction.
Figure 7: SLO optimization for BOT under LLAMA2-7B.
@@
(c) Avg TTFT and TPOT reduction.
Figure 8: SLO optimization for BOT under LLAMA2-13B.
@@
(c) Avg TTFT and TPOT reduction.
Figure 9: SLO optimization for SQL under LLAMA2-7B.
@@
(c) Avg TTFT and TPOT reduction.
Figure 10: SLO optimization for SQL under LLAMA2-13B.
Baselines. We compare SCOOT with the following baselines.
@@
  • Genetic Algorithm (GA): Mixed variable GA [30] is leveraged to reduce request tail latency and improve request throughput for CLS and REC applications, while NSGA2 [31] can handle multi-objective problems and is utilized to optimize TTFT and TPOT simultaneously for BOT and SQL applications. For GA, we assign a penalty objective value when constraints are violated. Besides, the population size of GA is set to ten.
  • Heteroscedastic Evolutionary BO (HEBO): HEBO [22] is a leading performance tuning framework that exploits Bayesian optimization to solve black-box optimization problems. HEBO can handle single-objective and multi-objective optimization problems with SOBO and MOBO, respectively. For HEBO, we assign a penalty objective value when constraints are violated.
(a) Tail latency reduction under LLAMA2-7B.
(b) Tail latency reduction under LLAMA2-13B.
(c) Avg tail latency reduction.
Figure 11: SLO optimization for CLS.
(a) Request TP improvement for LLAMA2-7B.
(b) Request TP improvement for LLAMA2-13B.
(c) Avg request TP improvement.
Figure 12: SLO optimization for REC.
V-B SLO Optimization
In the experiments of SLO optimization, we limit the total number of observations to 30 and set the suggestion parallelism degree (PD) for SCOOT to one. The performance of SLO optimization for SCOOT and baselines is presented in Figures 7, 8, 9, 10, 11, and 12. The bar represents the performance gain obtained by tuning, and the higher the bars, the better. In addition, × represents that no better configuration than the default configuration is found during the tuning process.
For BOT and SQL applications, TTFT and TPOT are optimized simultaneously. We present the tuning results of the two applications under different LLMs and computing resources, as well as the average tuning results across various computing resources in Figures 6, 7, 8, and 9. Parameter configurations obtained by tuning form a Pareto frontier of TTFT and TPOT, and we choose the configuration with the largest HV from the Pareto frontier to report since it represents the best trade-off between TTFT and TPOT. HV of a single configuration point $\boldsymbol{x}$ is calculated by
$HV(\boldsymbol{x})=\max(0,y_{\boldsymbol{x}}^{\phi}-y_{\boldsymbol{r}}^{\phi})\cdot\max(0,y_{\boldsymbol{x}}^{\theta}-y_{\boldsymbol{r}}^{\theta}),$ (11)
where $y_{\boldsymbol{x}}^{\phi}$, $y_{\boldsymbol{x}}^{\theta}$, $y_{\boldsymbol{r}}^{\phi}$, and $y_{\boldsymbol{r}}^{\theta}$ denote the observed TTFT and TPOT of $\boldsymbol{x}$ and the reference point $\boldsymbol{r}$ (i.e., the default configuration of the LLM inference engine), respectively.
From Fig 7, Fig. 8, Fig. 9, and Fig. 10, we can observe that for the average tuning performance in SLO optimization, SCOOT surpasses all the baselines in terms of both TTFT reduction and TPOT reduction under different LLMs, which confirms the SCOOT’s superiority in SLO optimization for interactive applications. For the BOT application, SCOOT decreases the TTFT and TPOT by up to 98.9% and 10.5% compared to the default configuration and baselines, respectively. In addition, for the SQL application, SCOOT reduces the TTFT and TPOT by up to 99.8% and 61.0% compared to the default configuration and baselines, separately.
While in some situations, SCOOT only outperforms the baselines on either TTFT or TPOT reduction, the optimized SLOs of SCOOT have a larger HV. For example, as shown in Fig. 8, in the case of 2A100, HEBO reduces TPOT slightly more than SCOOT, but SCOOT significantly outperforms HEBO in TTFT reduction, which leads to the optimized SLOs of SCOOT having an HV that is considerably larger than that of HEBO. Additionally, although in a very small number of cases, SCOOT failed to outperform the baseline in both TTFT and TPOT reductions, SCOOT’s tuning performance is more consistent. For example, as shown in Fig. 10, in the case of 2A100, HOBO outperforms SCOOT in both TTFT and TPOT reductions. However, in the case of 4A10, HEBO fails to find a configuration that is better than the default configuration, while SCOOT greatly reduces both TTFT and TPOT. In conclusion, SCOOT can always find superior parameter configurations. Such a consistently superior performance makes SCOOT outperform all the baselines in the average performance of SLO optimization across various computing resources.
For CLS and REC applications, request tail latency and request throughput are respectively utilized as optimization objectives. We present the tuning results of the two applications under different LLMs and computing resources in Figures 11 and 12, where we can observe that SCOOT always surpasses baselines in terms of tail latency reduction and throughput improvement under different LLMs, which confirms the SCOOT’s superiority in single-objective optimization.
For the CLS application, SCOOT reduces request tail latency by up to 40.6% compared to the default configuration and baselines. For the REC application, SCOOT decreases request throughput by up to 68.3% compared to the default configuration and baselines. Although HEBO’s performance is comparable to SCOOT and even slightly exceeds SCOOT in the case of 2A100 under LLAMA2-7B, as shown in Fig. 12, for LLAMA2-13B, HEBO’s performance is always far inferior to SCOOT, and hence SCOOT outperforms HEBO in average performance of SLO optimization for the REC application. This also confirms the SCOOT’s consistently superior performance on single-objective optimizations.
To sum up, for SLO optimization, SCOOT is always capable of finding superior parameter configurations for LLM inference services of various requests, LLMs, and computing resources to optimize various performance metrics.
(a) SLO optimization for BOT.
(b) SLO optimization for CLS.
Figure 13: Effectiveness of random forest-based POF learning.
SCOOT’s outstanding performance primarily comes from two aspects. First, SCOOT employs BO, which can efficiently handle black-box optimization problems with expensive evaluation overhead. Second, SCOOT can tackle complex constraints. By using hard constraints and learned hidden constraints to prune the solver’s solution space, it can significantly reduce the number of invalid observations without affecting the surrogate model. Additionally, this allows the surrogate model to fully learn the relationship between parameter configurations and optimization objectives. Therefore, SCOOT can always find good configurations within 30 observations.
GA optimizes around the initially sampled configurations. With a limited number of observations, both the population size and the number of iterations of GA are restricted to a very small value. Therefore, GA not only lacks diversity in solutions but also cannot fully utilize historical information to guide the search process, hence often performing poorly.
RD is unable to utilize historical information to guide the sampling process and can not handle constraints. Therefore, RD often fails to obtain sufficiently good configurations within 30 observations. However, due to the uniform sampling, RD can explore a large range in the search space and hence has a higher diversity in sampled configurations than GA. Hence, RD can often perform better than GA.
By leveraging BO to intelligently utilize historical data to guide the tuning process, HEBO often outperforms RD and GA in SLO optimization. However, HEBO simultaneously learns the objective function and constraints using the surrogate model. By assigning a penalty objective value to the unfeasible configurations, the learning of constraints hurts the modeling of the objective function. Configuration points near infeasible configurations are considered to be poorly performing points and are confused with values that are feasible but poorly performing. This can lead to a lack of subsequent exploration in the adjacent area of infeasible configurations, resulting in some potentially superior solutions being missed.
@@ V-B SLO Optimization
SCOOT can address this issue by leveraging GP as the surrogate model to approximate objective functions while learning hidden constraints with random forest, and hence the learning of hidden constraints does not affect the modeling of objective functions. In addition, by dynamically adjusting the POF threshold $\Delta$, SCOOT can explore the region near the infeasible configurations after continuous exploration of feasible configurations so as to explore the solution space more fully. We conduct experiments to confirm the effectiveness of the random forest-based learning of hidden constraints under 2A100 and 4A100. The results are presented in Fig. 13. Since requests from the CLS application seldom cause engine crashes, not applying RF-based POF learning has little impact on SCOOT’s performance. However, since BOT application requests cause many hidden constraints, RF-based POF learning can greatly enhance SLO optimization.
V-C Solution Quality

### v2

@@ 4.7. Random Forest-based POF Learning
Figure 5. Probability density function (PDF) of the request input and output lengths for four application request traces.
4.8. Ensuring SLO Robustness
@@
Through extensive experiments, we find that when the performance tuning is conducted under different request arrival orders, the best parameter configurations are almost the same, but the corresponding optimized SLOs vary. Thus, after performance tuning, we conduct stress testing multiple times again using the best configuration, varying request arrival orders, and select the worst-case objectives as the SLOs to ensure SLO Robustness. Finally, we summarize the overall tuning procedure of SCOOT in Algorithm 1.
5. Experimental Evaluation
5.1. Experiment Setup
We implement SCOOT on top of HEBO (cowen2022hebo, ), where the random forest regression is implemented using the sklearn (pedregosa2011scikit, ) library. Request traces are collected from four LLM inference services at Ant Group, including applications of text-to-SQL (SQL), chatbot (BOT), classification (CLS), and recommendation (REC). Since the computational load of an LLM inference request depends on its input and output lengths, to enhance the robustness of optimized SLOs, SCOOT uses requests with the longest 50% of output lengths for stress testing. The input and output lengths of these requests are presented in Fig. 5. For SQL and BOT, TTFT and TPOT are optimization objectives at the same time. For CLS and REC, 95th percentile latency and request throughput are utilized as optimization objectives, respectively.
@@ 5.1. Experiment Setup
Multiple sets of computing resources are utilized in experiments, including 2A10, 4A10, 2A100, and 4A100, where A10 represents the NVIDIA A10 24GB GPU, A100 denotes the NVIDIA A100 80GB GPU, and the integer before the GPU type indicate the number of GPUs. For each set of computing resources, 256GB CPU memory and one 2.90GHz 32-core Intel(R) Xeon(R) CPU are equipped. Besides, A100 GPUs are connected over NVLink while A10 GPUs are connected over PCIE. We use LLAMA2-7B and LLAMA2-13B (llama2, ) in experiments because the model architecture of LLAMA is representative and fine-tuned 7B and 13B LLMs can fulfill the requirements for most applications in practice. Due to the high usability and popularity of vLLM, we adopt vLLM-0.4.2 as the inference engine to tune and provide its known constraints in Appendix A.2.
@@
In each configuration evaluation, when the inference engine is started with the suggested configuration, we send requests from the request trace for 100 seconds and the request arrival times are generated using Poisson distribution with various request rates. For A10, the request rates for SQL, BOT, CLS, and REC are 5,5, 10, and 15, respectively. For A100, the request rates of these request traces are twice that of A10. The request rate is set according to the actual workload of the applications as well as considering the computing capability and GPU memory capacity of GPUs.
We compare SCOOT with three baselines, including random sampling (RD), genetic algorithm (GA), and Vanilla BO (VBO). Detailed baseline description is provided in Appendix A.3.
5.2. SLO Optimization
In the experiments of SLO optimization, we limit the total observation number to 30 and set the suggestion parallelism degree (PD) of SCOOT to one. The experimental results are presented in Figures 6, 7, 8, 9, 10, and 11, where we not only present the SLO improvement under each set of computing resources but also the average SLO improvement across various computing resources. The bar represents the SLO improvement compared to the default parameter configuration, and the higher the bars, the better. In addition, × represents that no better configuration than the default configuration is found.
For BOT and SQL applications, TTFT and TPOT are optimized simultaneously, and a Pareto frontier is resolved. We choose SLOs of the configuration with maximal HV from the Pareto frontier as the tuning result to report since it represents the best trade-off between TTFT and TPOT. HV of a single configuration point $\boldsymbol{x}$ is calculated by
(10) $HV(\boldsymbol{x})=\max(0,y_{\boldsymbol{x}}^{\phi}-y_{\boldsymbol{r}}^{\phi})\cdot\max(0,y_{\boldsymbol{x}}^{\theta}-y_{\boldsymbol{r}}^{\theta}),$
where $y_{\boldsymbol{x}}^{\phi}$, $y_{\boldsymbol{x}}^{\theta}$, $y_{\boldsymbol{r}}^{\phi}$, and $y_{\boldsymbol{r}}^{\theta}$ denote the observed TTFT and TPOT of $\boldsymbol{x}$ and the reference point $\boldsymbol{r}$, respectively.
@@ 5.2. SLO Optimization
Figure 6. SLO optimization for BOT under LLAMA2-7B.
@@
Figure 7. SLO optimization for BOT under LLAMA2-13B.
@@
Figure 8. SLO optimization for SQL under LLAMA2-7B.
@@
Figure 9. SLO optimization for SQL under LLAMA2-13B.
(a) LLAMA2-7B.
(b) LLAMA2-13B.
Figure 10. Tail latency reduction for CLS.
(a) LLAMA2-7B.
(b) LLAMA2-13B.
Figure 11. Request TP improvement for REC.
Figures 6, 7, 8, and 9 show that compared to the default configuration and baselines, SCOOT decreases the TTFT and TPOT by up to 98.9% and 10.5% for the BOT application, respectively, and reduces the TTFT and TPOT by up to 99.8% and 61.0% for the SQL application, separately. In addition, SCOOT surpasses all the baselines in terms of average TTFT and TPOT reduction across various computing resources for BOT and SQL applications, which confirms SCOOT’s superiority in multi-objective optimization.
Figures 10 and 11 present that compared to the default configuration and baselines, SCOOT reduces request tail latency by up to 40.6% for the CLS application and increases request throughput by up to 68.3% for the REC application. Besides, SCOOT outperforms all the baselines in terms of average tail latency reduction and request throughput improvement across various computing resources for CLS and REC applications, respectively, which confirms SCOOT’s superiority in single-objective optimization.
While in rare situations, SCOOT failed to outperform certain baselines in SLO optimization, SCOOT’s tuning performance is more consistent. As shown in Fig. 9, in the case of 2A100, VBO slightly outperforms SCOOT in both TTFT and TPOT reduction. However, in the case of 4A10, VBO fails to find a configuration that is better than the default configuration, while SCOOT greatly reduces TTFT and TPOT. Such a consistently superior performance makes SCOOT outperform all the baselines in the average performance of SLO optimization for various LLMs and applications.
GA optimizes around the initially sampled configurations. With a limited number of observations, both the population size and the number of iterations of GA are restricted to a small value. Thus, GA not only lacks diversity in solutions but also cannot fully utilize historical information to guide the search process, and GA can hardly tackle constraints, hence often performing poorly.
RD is unable to utilize historical information to guide the sampling process and can not handle constraints. However, due to the uniform sampling, RD can explore a large range in the search space and hence has a higher diversity in sampled configurations than GA. Hence, RD often performs better than GA.
By leveraging BO to intelligently utilize historical observations to guide the tuning process, VBO often outperforms RD and GA in SLO optimization. However, by assigning a penalty objective value to unfeasible configurations, the surrogate model will confuse points near infeasible configurations with points that are feasible but perform poorly. This can lead to a lack of subsequent exploration in the adjacent area of infeasible configurations, thus missing out on some potentially superior solutions.
SCOOT’s outstanding performance primarily comes from the employment of BO, which efficiently resolves black-box optimization problems, and the random forest-based POF learning, which intelligently handles hidden constraints without affecting the GP-based surrogate model. Besides, by dynamically adjusting the POF threshold $\Delta$, SCOOT can explore the region near the infeasible configurations after continuous exploration of feasible configurations so as to fully explore the solution space.
5.3. Tuning Efficiency
We conduct experiments to tune the CSL and BOT applications under 2A100 with different parallelism degrees (PDs). In the experiments, the number of observations is also limited to 30 when PD is set to 2. Figures 12(a) and 12(b) show the SCOOT’s performance in SLO optimization under various PDs, while Fig. 12(c) presents the total tuning time under various PDs.
Figures 12(a) and 12(b) show that when PD is set to 2, using the parallel suggestion achieves the same performance in SLO optimization as not using it. Besides, as presented in Fig. 12(c), when PD is set to 2, the parallel suggestion technique significantly reduces the total tuning time, nearly cutting it in half, which demonstrates that the parallel suggestion can effectively accelerate the tuning process by PD times. Furthermore, Figure 12(c) shows that compared with the time of configuration evaluation (i.e., stress testing), the time required for configuration suggestion is negligible, which validates the efficiency of applying BO to tune LLM inference engines.
We also increase PD and conduct experiments. However, we find that when PD is set to a large value, such as greater than 4, SCOOT’s performance of SLO optimization is compromised, and increasing the total observation number can mitigate this issue. In the future, we will explore the best number of observations under different PDs to find the most cost-effective acceleration solution.
(a) Tail latency reduction for CLS under LLAMA2-7B & 13B.
(b) TTFT & TPOT reduction for BOT under LLAMA2-7B.
(c) Performance tuning time for CLS and BOT.
Figure 12. Tuning efficiency of the parallel suggestion.
5.4. Supplementary Experiments
In the appendix, we evaluate the effectiveness of the random forest (RF)-based learning of hidden constraints through ablation studies, confirm the superiority of SCOOT in both the optimality and diversity of SLO optimization in multi-objective situations, and show that SCOOT can outperform baselines and human experts through real-world inference service optimization results, where both vLLM and TensorRT-LLM are leveraged to verify the universality of SCOOT.
6. Related Works
6.1. Advanced LLM Inference Techniques
Existing inference engines such as vLLM and TensorRT-LLM support a variety of advanced inference techniques to improve the speed and throughput of LLM inference, including a series of kernel-level optimizations such as continuous batching (yu2022orca, ), paged attention (kwon2023efficient, ), flash attention (dao2022flashattention, ; dao2023flashattention2, ), and chunked prefill (agrawal2024taming, ). Besides, prefix caching (jhaimproving, ) is employed by inference engines to reuse computed key and value tensors of requests’ common prefix texts to reduce redundant memory allocation and computation for newly arrived requests, which can significantly improve the serving efficiency when requests share a long common prefix. Moreover, speculative decoding (leviathan2023fast, ; xia2023speculative, ; miao2024specinfer, ; kim2024speculative, ; sun2024spectr, ) is also adopted by inference engines to accelerate LLM inference, where an efficient draft model is leveraged to generate tokens, and the LLM occasionally refines the draft.
6.2. Automatic Performance Tuning Approaches
Performance tuning techniques have been widely applied across a wide variety of fields, including database tuning (zhang2021restune, ; lao2023gptuner, ; yang2024vdtuner, ), Spark configuration tuning (gounaris2017dynamic, ; fekry2020tune, ; li2023towards, ; shen2023rover, ), compiler optimization (hellsten2023baco, ), and tuning of web-relevant applications (dalibard2017boat, ; li2018bayesian, ). The vast majority of these performance tuning studies adopt BO to find optimized parameter configurations since BO is theoretically grounded and can efficiently learn the relationship between performance and parameters from evaluations, thus intelligently tuning parameters for performance improvement. Because of these advantages, BO is also employed in resource allocation studies (patel2020clite, ; roy2021satori, ; yang2023cotuner, ).
To the best of our knowledge, previous studies have not focused on tuning LLM inference engines and can not address all the three unique challenges faced in inference engine tuning.
7. Conclusion
In this paper, we propose SCOOT to optimize SLOs for LLM inference services via BO-based automatic inference engine tuning. SCOOT uses RF to learn hidden constraints and employs the parallel suggestion to speed up tuning. Extensive experiments are conducted to confirm both the effectiveness and efficiency of SCOOT.
Acknowledgements.
This work is supported in part by the Postgraduate Research & Practice Innovation Program of Jiangsu Province (KYCX24_0247), the Nanjing Key S&T Special Projects (202309006), and the Ant Research Program of Ant Group.
Appendix A Appendix
A.1. Conflicting TTFT and TPOT
(a) TTFT-only optimization.
(b) TPOT-only optimization.
Figure 13. Conflicting TTFT and TOPT. Decreased TTFT often leads to increased TPOT, and vice versa. The red line indicates the objective value under the default parameter configuration.
(a) Tail latency reduction for CLS under LLAMA2-7B & 13B.
(b) TTFT & TPOT reduction for BOT under LLAMA2-7B.
Figure 14. Effectiveness of random forest-based POF learning.
We separately take TTFT and TPOT as the only optimization objectives and conduct grid searches for inference services of LLAMA2-13B deployed on various computing resources under the BOT request trace. Experimental results are shown in Fig. 13, which confirms that for TTFT and TPOT, optimizing for one always compromised the other. The reason is that, to decrease TPOT, it is significant to decrease the batch size controlled by the parameter max-num-seqs to reduce the per-token latency, which causes a long queuing time for requests, thus leading to a large TTFT. Conversely, to reduce TTFT, a large batch size is always set to allow newly arrived requests to be processed promptly without queuing for a long time. However, a large batch size often slows down inference, leading to increased TPOT.
A.2. Known Constraints of Inference Engines
A.2.1. vLLM
For vLLM-0.4.2 and vLLM-0.5.5, the known constraints include that max-num-batched-tokens must be greater than or equal to max_num_seqs, the enable-chunked-prefill and enable-prefix-caching cannot be set to True simultaneously, and when the enable-chunked-prefill is not enabled, max-num-batched-tokens needs to be greater than or equal to the max model length specified in the model configuration file.
A.2.2. TensorRT-LLM
For TensorRT-LLM-0.15.5, the known constraints include that max-num-tokens must be greater than or equal to max_batch_size, use-paged-context-fmha must be set to True if enable-chunked-context is enabled, and when the enable-chunked-contex is not enabled, max-num-tokens needs to be greater than or equal to the max model length specified in the model configuration file.
A.3. Baseline Description
@@
  • Genetic Algorithm (GA): Mixed variable GA [47] is used to resolve single-objective optimization problems while NSGA2 [48] is utilized to handle multi-objective situations. For GA, we assign a penalty objective value when constraints are violated and set the population size to 10.
  • Vanilla BO (VBO): VBO uses UCB and EHVI as acquisition functions to handle single-objective and multi-objective optimization situations, respectively. For VBO, we assign a penalty objective value when constraints are violated and use a random forest-based surrogate model to learn objective functions and constraints at the same time.
All of these approaches are developed using HEBO [22], which has highly optimized implementations. Therefore, we think they are competitive and suitable as baselines.
A.4. Ablation Study Experiments
@@
We conduct experiments to confirm the effectiveness of the random forest (RF)-based learning of hidden constraints under 2A100 and 4A100. The average SLO optimization results across various computing resources are presented in Fig. 14. Since requests from the CLS application seldom cause engine crashes, not applying RF-based POF learning has little impact on SCOOT’s performance. However, since BOT application requests cause many hidden constraints, RF-based POF learning can greatly enhance SLO optimization.
(a) Optimized TTFT & TPOT under LLAMA2-7B.
(b) Optimized TTFT & TPOT under LLAMA2-13B.
Figure 15. Solution quality for BOT.
(a) Optimized TTFT & TPOT under LLAMA2-7B.
(b) Optimized TTFT & TPOT under LLAMA2-13B.
Figure 16. Solution quality for SQL.
A.5. Solution Quality Experiments